OpenStudy (anonymous):

Solve the equation -4x+7y+7=0 for y

OpenStudy (anonymous):

-4x+7y+7=0 -4x+7y=-7 7y=4x-7 y=4/7x-1

OpenStudy (anonymous):

correct ^_^